OpenStudy (ammarah):

Solve each equation by completing the square. if necessary found to the neares hundred.

OpenStudy (ammarah):

x^2-x=2

OpenStudy (ammarah):

@ganeshie8

ganeshie8 (ganeshie8):

take half of "x" coefficient

ganeshie8 (ganeshie8):

add its square both sides

OpenStudy (ammarah):

ok so 1^2

ganeshie8 (ganeshie8):

what is "x" coefficient ?

OpenStudy (ammarah):

1

ganeshie8 (ganeshie8):

whats half of 1 ?

OpenStudy (ammarah):

1/2

ganeshie8 (ganeshie8):

\(\large x^2-x=2\) \(\large x^2-2\times \frac{1}{2}x=2\)

ganeshie8 (ganeshie8):

so, add the missing \(\left(\frac{1}{2}\right)^2\) both sides, so that we can comlete the square

OpenStudy (ammarah):

why is i t times

ganeshie8 (ganeshie8):

\(\large x^2-x=2 \) \(\large x^2-2\times \frac{1}{2}x=2\) \(\large x^2-2\times \frac{1}{2}x + \left(\frac{1}{2}\right)^2=2 + \left(\frac{1}{2}\right)^2\)

ganeshie8 (ganeshie8):

In the second step, I have just wrote 1 as 2*1/2

ganeshie8 (ganeshie8):

2*1/2 = 1, so... nothing fishy ok ?

ganeshie8 (ganeshie8):

Next, look at the left hand side once, does the pattern look familiar ?

ganeshie8 (ganeshie8):

\(\large a^2 - 2a b + b^2 \)

ganeshie8 (ganeshie8):

??

OpenStudy (ammarah):

what

ganeshie8 (ganeshie8):

you should know this identity : \(\large a^2 -2ab + b^2 = (a-b)^2 \)

ganeshie8 (ganeshie8):

the left hand side is exactly of above form, so we can write it as a perfect square

ganeshie8 (ganeshie8):

\(\large x^2-x=2 \) \(\large x^2-2\times \frac{1}{2}x=2\) \(\large x^2-2\times \frac{1}{2}x + \left(\frac{1}{2}\right)^2=2 + \left(\frac{1}{2}\right)^2\) rewriting the left hand side using a known identity : \(a^2-2ab+b^2 = (a-b)^2\) \(\large \left( x - \frac{1}{2}\right)^2 =2 + \left(\frac{1}{2}\right)^2\) simplifying the right hand side : \(\large \left( x - \frac{1}{2}\right)^2 = \frac{9}{4}\)

ganeshie8 (ganeshie8):

take square root both sides

ganeshie8 (ganeshie8):

\(\large x^2-x=2 \) \(\large x^2-2\times \frac{1}{2}x=2\) \(\large x^2-2\times \frac{1}{2}x + \left(\frac{1}{2}\right)^2=2 + \left(\frac{1}{2}\right)^2\) rewriting the left hand side using a known identity : \(a^2-2ab+b^2 = (a-b)^2\) \(\large \left( x - \frac{1}{2}\right)^2 =2 + \left(\frac{1}{2}\right)^2\) simplifying the right hand side : \(\large \left( x - \frac{1}{2}\right)^2 = \frac{9}{4}\) take square root both sides : \(\large x - \frac{1}{2} = \pm \frac{3}{2}\)

ganeshie8 (ganeshie8):

you get : \(\large x = 2, -1\)

ganeshie8 (ganeshie8):

we're done. let me knw if somthng doesnt make sense